A big lump of meat of mass 5kg is hung from a spring balance in an elevator , find the reading of the balance if
1) the elevator is moving with a steady speed
2)the elevator is moving upwards with acceleration of 0.2ms-
1
Expert's answer
2021-08-23T12:41:33-0400

The reading of the balance is actually the weight of the lump.

1. In this case the reference frame connected to the elevator can be considered as the intertial one (no extra acceleration). Thus, the weight of the lump is the same as its weight in the stationary lift:


"P = mg = 5kg\cdot 9.8m/s^2 = 49N"

2. In this case the weight is given as follows (see https://physicsnoteonline.blogspot.com/p/blog-page_29.html):


"P = m(g +a) = 5kg\cdot (9.8m/s^2 + 0.2m/s^2) = 50N"

Answer. 1) 49N, 2) 50N.